# Gromov-Hausdorff limits of smocked spaces

### v2

@@ 3.2 Subsequential Gromov-Hausdorff convergence
In [2], it was shown that sequences of re-scalings of smocked spaces do converge in the pointed Gromov-Hausdorff sense .even though the spaces themselves do not satisfy the hypotheses of the Gromov pre-compactness theorem. There are some immediate examples of Gromov-Hausdorff converging subsequences which one can construct. For example, take a sequence of smocked spaces $\{X_{k}\}_{k=1}^{\infty}$ such that $X_{k}$ is a smocked space with one single smocking stitch for odd $k$ and such that $X_{k}$ is a smocked space with two smocking stitches which are the same compact subsets $\mathbb{E}^{n}$ for each even $k$. This sequence of smocked spaces contains a Gromov-Hausdorff converging subsequence given by $(X_{2k})$.
Generally speaking, we would like conditions for a sequence of smocked spaces to contain a pointed Gromov-Hausdorff converging subsequence. It turns out that to get such conditions it is sufficient to assume uniform bounds on the smocking constants. Intuitively, the result says that one cannot pack in too many sets too closely together in the smocked space. This may explain why the examples found by Sormani et al. can be shown to have rescaled sequences which converge even though smocked spaces themselves do not meet the hypotheses of Gromov’s compactness theorem. As stated earlier, one obtains a sequence of metric spaces when a metric space is repeatedly rescaled and the Gromov-Hausdorff limit (if it exists) is the tangent cone at infinity.
Proposition 3.2: A sequence of smocked spaces $\{X_{k}\}_{k=1}^{\infty}$ contains a Gromov-Hausdorff converging subsequence if there are uniform upper bounds on the smocking constants.
Proof: The proof is similar to others in the literature, but we will reproduce the details here for completeness [1]. We must prove that there is a finite number of indices $i_{1},...,i_{p}$ such that for every $k$, there exists $j$ with
$d_{GH}((B_{r}(x_{0}^{(k)}),dX_{k}),(B_{r}(x_{0}^{(i_{j})}),dX_{i_{j}}))\leq\text{constant}\cdot D_{k},$
where $D_{k}$ is defined to be $\text{sup}\{L_{\text{min}}^{k},L_{\text{max}}^{k},h_{k},\delta_{k}\}$. $D_{k}$ is assumed to be finite by hypothesis.
Since $D_{k}$ is finite, for each ball $(B_{r}(x_{0}^{(k)}),dX_{k})$ the maximal disjoint filling is achieved by balls of radius $cD_{k}$ centred at points of a $2cD_{k}$-net of the ball (call the net $L$), for some $c>0$. The number of balls in the filling is bounded by some quantity $N=N(D_{k},x_{0}^{(k)},r)$. A choice of bijection $f:\{1,...,N\}\longrightarrow L$ induces a distance function $d_{L}$ on the set $\{1,...,N\}$. The function $d_{L}:\{1,...,N\}\times\{1,...,N\}\longrightarrow\mathbb{R}_{+}$ which assumes values between $2cD_{k}$ and $2r$ on the subset of $S$ of pairs of different integers has $N^{2}-N$ elements. The space $[2cD_{k},2r]^{N^{2}-N}$ is precompact with respect to the product distance function, so there exist finitely many indices $i_{1},...,i_{p}$ such that, for each $i$, there exists $j$ such that
$\sup_{S}|d_{k}(m,q)-d_{i_{j}}(m,q)|\leq c^{2}D_{k}^{2},$
which implies that
$\sup_{S}\frac{d_{k}(m,q)}{d_{i_{j}}(m,q)}\leq\frac{1+cD_{k}}{1-cD_{k}}.$
This in turn implies that
$d_{H}((B_{r}(x_{0}^{(k)}),dX_{k}),(B_{r}(x_{0}^{(i_{j})}),dX_{i_{j}}))\leq(2r+4)cD_{k}.$
Finally, Hausdorff convergence here implies pointed Gromov-Hausdorff convergence, so we have
$d_{GH}((B_{r}(x_{0}^{(k)}),dX_{k}),(B_{r}(x_{0}^{(i_{j})}),dX_{i_{j}}))\leq(2r+4)cD_{k},$
@@
Remark 3.3: Note that having uniform bounds on the smocking constants is a sufficient, not a necessary, condition for subsequential convergence. To show that it is not necessary, one may consider a sequence where each space $X_{k}$ consists of two ’stitches’ which move further apart with increasing $k$. This converges to $\mathbb{E}^{N}$ in the limit, but the smocking separation factor $\delta$ is unbounded.
3.3 Normed spaces from smocked spaces
@@
Finally, we show that one can always obtain a finite-dimensional normed vector space as a pointed Gromov-Hausdorff limit of a sequence of smocked spaces. In general, there are examples of smocked spaces which do not have unique tangent cones at infinity, nor is the cone always a normed space [2].
Proposition 3.4: Given a finite-dimensional normed vector space $(Y,d_{Y})$, there exists a smocked space $(X,d_{X})$ whose tangent cone at infinity is $(Y,d_{Y})$.
Proof: For a metric space $(X,d)$ to be a normed space, the metric $d$ must be translation invariant
$d(x,y)=d(x+a,y+a),\>\>\>\>\>\>\>\forall a\in X$
and homogeneous
$d(cx,cy)=cd(x,y),\>\>\>\>\>\>\>\forall c\in\mathbb{R}.$
@@ 4.1 Measured convergence of smocked measure spaces
A natural question is whether the results which were derived above still apply for metric spaces equipped with measures. We will begin by recalling the definition of metric measure spaces given in [7].
Definition 4.1 A metric measure space is a triple $(X,d,\mu)$, where $(X,d)$ is a metric space and $\mu$ is a measure on the Borel $\sigma$-algebra of $X$.
We then propose to extend this definition to smocked spaces in order to work with what we will call pointed smocked metric measure spaces (or smocked measure spaces for short). As before, we must use pointed convergence because smocked spaces are non-compact.
Definition 4.2 A smocked metric measure space is a triple $(X,d,\mu)$, where $(X,d)$ is a smocked space and $\mu$ is a non-zero Borel measure which is finite on bounded sets. A pointed smocked metric measure space $(X,d,\mu,\overline{x})$ is a smocked metric measured space together with a given point $\overline{x}\in\text{supp}(\mu)\subset X$.
The first question we might ask is whether we can obtain analogues of our previous convergence results established above, but now for the case of smocked measure spaces. We will begin by defining a suitable notion of pointed measured Gromov-Hausdorff convergence used by Lott and Villani and adapted by Gigli, Mondino and Savaré [8-10]. This notion works with proper (that is, bounded closed subsets are compact) pointed metric measure spaces. $\mathbb{E}^{N}$ is proper finite-dimensional, so distances converge uniformly on bounded sets.
Definition 4.3 If one has a sequence of pointed smocked metric measure spaces $(X_{j},d_{j},\mu_{j},\overline{x}_{j})$, pointed measured Gromov-Hausdorff convergence is defined as follows:
$(X_{j},d_{j},\mu_{j},\overline{x}_{j})\xrightarrow{pmGH}(X_{\infty},d_{\infty},\mu_{\infty},\overline{x}_{\infty})$
if and only if for every $\epsilon,R>0$ there exists $N(\epsilon,R)$ such that, for all $n\geq N$, there exists a Borel map $f_{j}^{R,\epsilon}$: $B_{R}(\overline{x}_{j})\rightarrow\overline{x}_{\infty}$ such that
(i) $f_{j}^{R,\epsilon}(\overline{x}_{j})=\overline{x}_{\infty}$;
(ii) $\text{sup}_{x,y\in B_{R}(\overline{x}_{j})}|d_{j}(x,y)-d_{\infty}(f_{j}^{R,\epsilon}(x),(f_{j}^{R,\epsilon}(y))|\leq\epsilon$;
(iii) the $\epsilon$-neighbourhood of $f_{j}^{R,\epsilon}(B_{R}(\overline{x}_{j}))$ contains $B_{R-\epsilon}(\overline{x}_{\infty})$;
(iv) $(f_{j}^{R,\epsilon})_{\#}(\mu_{j}\rvert_{B_{R}(\overline{x}_{j})})$ weakly converges to $\mu_{\infty}\rvert_{B_{R}(\overline{x}_{\infty})}$ as $n\rightarrow\infty$, for almost every $R>0$.
This definition is somewhat more complicated than the one for pointed Gromov-Hausdorff convergence, but we can still show that if one has a sequence of smocking sets $S_{k}$ which Hausdorff converge to $S_{\infty}$ as before, then we can always define corresponding smocked measure spaces $(X_{k},\mu_{K})$ which converge to a pointed measured Gromov-Hausdorff limit. By Proposition 3.1, it is only necessary to show that pointed Gromov-Hausdorff convergence implies pointed measured Gromov-Hausdorff convergence.
Proposition 4.4: If a sequence of smocked spaces $(X_{k},x_{k},d_{k})$ converges in the pointed Gromov-Hausdorff sense to $(X_{\infty},x_{\infty},d_{\infty})$, then one can define a sequence of smocked measure spaces $(X_{k},d_{k},\mu_{k},\overline{x}_{k})$ which converges in the pointed measured Gromov-Hausdorff sense to $(X_{\infty},d_{\infty},\mu_{\infty},\overline{x}_{\infty})$.
@@
The following corollary is immediate.
Corollary 4.5: If one defines a sequence of smocked measure spaces using smocking sets $S_{k}\subset\mathbb{E}^{N}$ such that the sets Hausdorff converge in the limit to a smocking set $S_{\infty}$, then one can define corresponding smocked measure spaces which converge to the corresponding limit space in the pointed measured Gromov-Hausdorff sense.
We conjecture that it should be possible to show sufficient conditions for subsequential pointed measured Gromov-Hausdorff convergence of smocked measure spaces by explicitly constructing an abstract limit space with an ancillary measure and showing that convergence is obtained after extraction of subsequences.
Conjecture 4.6: A sequence of smocked measure spaces $\{(X_{k},d_{k},\mu_{k},\overline{x}_{k})\}_{k=1}^{\infty}$ contains a subsequence which converges in the pointed measured Gromov-Hausdorff sense if there are uniform upper bounds on both the smocking constants and $\text{sup}\>\mu_{k}(B_{r}(x_{0}^{(k)}))$.
Remark 4.7: Much recent work has focussed on stability of metric measure spaces under different types of convergence. As an example, one may consider stability of the curvature dimension condition under pointed measured Gromov-Hausdorff convergence [13, 14, 15-17,10]. There is no such stability result for smocked measure spaces because they are not length spaces nor are they compact, which means that the associated Wasserstein spaces also cannot be compact or length spaces.
4.2 Smocked Wasserstein Spaces
Pulled thread spaces (and more generally, smocked spaces) can be defined by starting with any geodesic metric space or length space [1]. We conjecture here that Wasserstein spaces which are also length spaces can be smocked by considering smocking sets which are disjoint connected compact subset of the space of Borel probability measures. It is possible that smocking the Wasserstein spaces would have an influence on $GW_{p}$-convergence of the original metric measure spaces.
Another basic principle of optimal transport is that a pair of Borel probability measures on a manifold which evolve under a diffusion equation should converge in the Wasserstein distance if the manifold satisfies a curvature condition (usually positive Ricci curvature as defined above) [19]. The work of Sturm and von Renesse shows that if a smooth, compact Riemannian manifold $M$ is equipped with two Borel probability measures $\mu$ and $\nu$ which spread according to the heat equation, then the lower bound on the Ricci curvature of $M$ controls the speed with which the Wasserstein distance between the two measures can expand or decay over time. In particular, it was shown that the following properties are all equivalent for such a manifold:
Proposition 4.8 [20]: For $K\in\mathbb{R}$ and a compact Riemannian manifold $M$, the following are equivalent:
(i) $\text{Ric}(M)\geq K$.
(ii) For all $x$ and $y$ in $M$ and all $t>0$, there exists $r\in[1,\infty]$ with
$d_{W_{P}}(p_{t}(x,\cdot),p_{t}(y,\cdot))\leq e^{-Kt}\cdot d(x,y).$
(iii) For all $r\in[1.\infty]$, all $\mu$ and $\nu$ in $\mathcal{P}_{P}(M)$, and all $t>0$
$d_{W_{P}}(\mu p_{t}(x,\cdot),\nu p_{t}(y,\cdot))\leq e^{-Kt}\cdot d_{W_{P}}(\mu,\nu),$
where $p_{t}(x,y)$ denotes the minimal positive fundamental solution of the heat equation.
We conjecture that if a smocked Wasserstein distance can be defined, condition (iii) can be replaced with the same statement but using this new distance.

### v3

@@ 3.2 Subsequential Gromov-Hausdorff convergence
In [2], it was shown that sequences of re-scalings of smocked spaces do converge in the pointed Gromov-Hausdorff sense, even though the spaces themselves do not satisfy the hypotheses of the Gromov pre-compactness theorem. There are some immediate examples of Gromov-Hausdorff converging subsequences which one can construct. For example, take a sequence of smocked spaces $\{X_{k}\}_{k=1}^{\infty}$ such that $X_{k}$ is a smocked space with one single smocking stitch for odd $k$ and such that $X_{k}$ is a smocked space with two smocking stitches which are the same compact subsets $\mathbb{E}^{n}$ for each even $k$. This sequence of smocked spaces contains a Gromov-Hausdorff converging subsequence given by $(X_{2k})$.
Generally speaking, we would like conditions for a sequence of smocked spaces to contain a pointed Gromov-Hausdorff converging subsequence. It turns out that to get such conditions it is sufficient to assume uniform global bounds on the smocking constants. Intuitively, the result says that one cannot pack in too many sets too closely together in the smocked space. This may explain why the examples found by Sormani et al. can be shown to have rescaled sequences which converge even though smocked spaces themselves do not meet the hypotheses of Gromov’s compactness theorem. However, to also have necessity, one must assume stronger local bounds on the smocking constants and local geometric control.
We begin by recalling Gromov’s precompactness theorem.
Proposition 3.4: Let $\{(Y_{\alpha},y_{\alpha})\}$ be a family of pointed proper metric spaces (i.e. all closed metric balls are compact). If for every $r,\varepsilon>0$ there is $N(r,\varepsilon)<\infty$ such that each $B_{r}^{Y_{\alpha}}(y_{\alpha})$ can be covered by at most $N(r,\varepsilon)$ balls of radius $\varepsilon$, then the family is precompact in the pointed Gromov–Hausdorff topology.
@@
We next show that global bounds on the smocking constants are sufficient for Gromov-Hausdorff precompactness.
Proposition 3.5: Let $\{X_{k}\}$ be a sequence of smocked spaces obtained from $\mathbb{E}^{N}$ by collapsing collections of compact stitches, and let $\pi_{k}:\mathbb{E}^{N}\to X_{k}$ be the smocking maps. Fix basepoints $x_{k}=\pi_{k}(0)$ and assume that there exists a set of smocking constants $L_{\max}<\infty$, $h<\infty$, $\delta>0$ and $L_{\min}>0$ independent of $k$ such that $L^{(k)}_{\text{max}}\leq L_{\text{max}}$, $h^{(k)}\leq h$, $\delta^{(k)}\geq\delta$, and $L^{(k)}_{\text{min}}\geq L_{\text{min}}$. Then the family $\{(X_{k},x_{k})\}$ is precompact in the pointed Gromov-Hausdorff topology (every sequence admits a subsequence which converges in the pointed Gromov-Hausdorff sense).
Proof: Fix $r>0$ and $\varepsilon>0$. The strategy of the proof is to produce a uniform bound $N(r,\varepsilon)$ on the number of disjoint $\varepsilon$-balls that can be packed into any ball $B_{r}^{X_{k}}(x_{k})$, with the required subsequence then following from Gromov’s precompactness theorem. By properties of the smocking map, we have for every $k$
$\pi_{k}^{-1}\!\big(B_{r}^{X_{k}}(x_{k})\big)\subset B_{r+L_{\max}}^{\mathbb{E}^{N}}(0).$
Since distinct stitches are separated by at least $\delta>0$ and every stitch has diameter $\leq L_{\max}$, only finitely many stitches intersect the compact Euclidean ball $B_{r+L_{\max}}^{\mathbb{E}^{N}}(0)$, with a bound which depends only on $r,L_{\max},\delta$ and $N$.
Next, set
$\varepsilon^{\prime}:=\min\{\varepsilon/4,\ \delta/4\}.$
Suppose that $\{B_{\varepsilon}^{X_{k}}(y_{i})\}_{i=1}^{M}$ is a collection of pairwise disjoint $\varepsilon$–balls contained in $B_{r}^{X_{k}}(x_{k})$. Choose lifts $v_{i}\in\pi_{k}^{-1}(y_{i})\subset B_{r+L_{\max}}^{\mathbb{E}^{N}}(0)$. We next claim that the Euclidean distance between any two distinct lifts satisfies $|v_{i}-v_{j}|\geq\varepsilon^{\prime}$. To see this, consider any two distinct points $y_{i},y_{j}\in X_{k}$ with $d_{X_{k}}(y_{i},y_{j})\geq\varepsilon$, and choose lifts $v_{i}\in\pi_{k}^{-1}(y_{i})$, $v_{j}\in\pi_{k}^{-1}(y_{j})$. There are two cases which must be checked separately.
In the first case, the lifts $v_{i},v_{j}$ lie in the same component of $\mathbb{E}^{N}\setminus\!\bigcup_{\alpha}S_{\alpha}^{(k)}$. $\pi_{k}$ is an isometry on that component, so we have $|v_{i}-v_{j}|_{\mathbb{E}^{N}}=d_{X_{k}}(y_{i},y_{j})\geq\varepsilon$. In the second case, the lifts lie in different components, hence they are separated by at least one collapsed stitch. Distinct stitches are $\delta$-separated in $\mathbb{E}^{N}$, so $|v_{i}-v_{j}|\geq\delta$. In particular, no two lifts of distinct $\varepsilon$–separated points in $X_{k}$ can be closer than $\varepsilon^{\prime}:=\min\{\varepsilon/4,\delta/4\}$ in $\mathbb{E}^{N}$. In both cases, we obtain $|v_{i}-v_{j}|\geq\varepsilon^{\prime}$ as claimed. It follows that the Euclidean balls $B_{\varepsilon^{\prime}/2}^{\mathbb{E}^{N}}(v_{i})$ are pairwise disjoint and lie inside the fixed compact set $B_{r+L_{\max}+\varepsilon^{\prime}/2}^{\mathbb{E}^{N}}(0)$.
By the Euclidean volume packing estimate,
$M\leq\frac{\operatorname{Vol}\big(B_{r+L_{\max}+\varepsilon^{\prime}/2}(0)\big)}{\operatorname{Vol}\big(B_{\varepsilon^{\prime}/2}(0)\big)}=:N(r,\varepsilon),$
which depends only on $r,\varepsilon,L_{\max},\delta$ and $N$ (in particular it is independent of $k$). As a result, the number of disjoint $\varepsilon$–balls that can be packed into $B_{r}^{X_{k}}(x_{k})$ is uniformly bounded by $N(r,\varepsilon)$.
[…]
Remark 3.6: Note that having uniform bounds on the smocking constants is a sufficient, but not a necessary, condition for subsequential convergence. For clarity, we will provide a counterexample in one dimension, but the argument can be easily generalized. For each integer $k\geq 1$, let $S_{k}=[k^{2},k^{2}+k]\subset\mathbb{R}$ and let $X_{k}$ be the smocked space obtained from $\mathbb{R}$ by collapsing $S_{k}$ to a single point. Denote by $\pi_{k}:\mathbb{R}\to X_{k}$ the smocking map and take the basepoint $x_{k}=\pi_{k}(0)$.
The maximal stitch diameter satisfies
$L_{\max}^{(k)}=\operatorname{diam}(S_{k})=k\longrightarrow\infty,$
so there is no global upper bound on $L_{\max}^{(k)}$. However, for any fixed radius $R>0$ the collapsed interval $S_{k}$ lies outside $\pi_{k}^{-1}\!\big(B_{R}^{X_{k}}(x_{k})\big)$ once $k$ is large, therefore the restriction of the metric on $B_{R}^{X_{k}}(x_{k})$ agrees with the Euclidean metric on [R,R][-R,R]. As a result, we have
$(X_{k},x_{k})\xrightarrow{\mathrm{pGH}}(\mathbb{R},0).$
This example shows that pointed Gromov-Hausdorff precompactness does not require uniform bounds on the smocking constants.
The assumption of global uniform bounds on the smocking constants is rather strong, since it controls the smocking pattern for all $k$ and across the entire space. In reality, pointed Gromov-Hausdorff convergence only sees geometry in neighborhoods around basepoints, so one may suspect that having bounded smocking constants should be both necessary and sufficient for pointed Gromov-Hausdorff convergence if the bounds are controlled in some local sense. The next proposition will focus on making this intuition precise.
Proposition 3.7: Let $\{(X_{k},x_{k})\}$ be a sequence of smocked spaces obtained from $\mathbb{E}^{N}$ by collapsing collections of compact stitches, with smocking maps $\pi_{k}:\mathbb{E}^{N}\to X_{k}$ and basepoints $x_{k}=\pi_{k}(0)$. The sequence $\{(X_{k},x_{k})\}$ admits a subsequence which converges in the pointed Gromov–Hausdorff sense if and only if the following local uniform control holds: for every $r>0$, there exist smocking constants $L_{r}<\infty$, $\delta_{r}>0$ and an index $K_{r}$ such that for all $k\geq K_{r}$, every stitch that intersects $\overline{B_{r+L_{r}}^{\mathbb{E}^{N}}(0)}\;$ has diameter less than or equal to $L_{r}$ and any two distinct such stitches are separated by Euclidean distance at least $\delta_{r}$.
Proof: Begin by assuming that $\{(X_{k},x_{k})\}$ has a subsequence which converges in the pointed Gromov-Hausdorff sense. For simplicity, relabel and assume the whole sequence converges to a pointed limit $(X_{\infty},x_{\infty})$ (the argument below applies to any convergent subsequence). Fix $r>0$. By pointed Gromov-Hausdorff convergence, we have for every $\varepsilon>0$ a uniform packing bound $N(r,\varepsilon)$ for the balls $B_{r}^{X_{k}}(x_{k})$ for all sufficiently large $k$. Concretely, choose $\varepsilon<1$ and let $M:=N(r,\varepsilon)$. If for some large $k$ there were infinitely many distinct stitches intersecting $\overline{B_{r+1}^{\mathbb{E}^{N}}(0)}$, then by picking one point in each stitch and projecting to $X_{k}$ we would obtain arbitrarily many pairwise $\varepsilon$-separated points in $B_{r}^{X_{k}}(x_{k})$, contradicting the packing bound. It follows that for all large $k$ only finitely many stitches meet $\overline{B_{r+1}(0)}$, so the diameters of those stitches are uniformly bounded from above and their pairwise separations are uniformly bounded from below by positive constants. This gives the required $L_{r}$ and $\delta_{r}$ and an index $K_{r}$, so that the stated inequalities hold for all $k\geq K_{r}$.
For the other direction of the implication, begin by assuming that the local bounds exist for every $r>0$. Fix arbitrary $r>0$ and $\varepsilon>0$. By hypothesis there is $L_{r},\delta_{r}$ and an index $K_{r}$ such that for all $k\geq K_{r}$ every stitch meeting $\overline{B_{r+L_{r}}(0)}$ has diameter $\leq L_{r}$ and distinct such stitches are $\geq\delta_{r}$ apart. The argument in Proposition 3.3 can now be applied on the fixed radius $r$. Lift any collection of pairwise-disjoint $\varepsilon$-balls in $B_{r}^{X_{k}}(x_{k})$ to $\mathbb{E}^{N}$ and use the uniform separation and diameter bounds to produce a uniform Euclidean packing of radius $\varepsilon^{\prime}=\min\{\varepsilon/4,\delta_{r}/4\}$ in the fixed compact set $B_{r+L_{r}+\varepsilon^{\prime}/2}(0)$. It follows that there is a uniform bound $N(r,\varepsilon)$ on the number of disjoint $\varepsilon$-balls in $B_{r}^{X_{k}}(x_{k})$, valid for all $k\geq K_{r}$.
[…]
The previous proposition shows that that there is a geometric criterion for precompactness in terms of the smocking pattern: the constants $L_{r}$ and $\delta_{r}$ control, respectively, the maximal size of a stitch and the minimal Euclidean separation between distinct stitches inside a ball $B_{r+L_{r}}^{\mathbb{E}^{N}}(0)$. The smocking depth $h$ controls the geometry, but does not play a role in convergence and does not appear in the assumptions used for the proofs. The local bound hypothesis requires that smocking stitches should remain under uniform control within fixed bounded regions around the basepoint, although the geometry of the stitches may degenerate at arbitrarily large distances (stitches whose diameters tend to infinity whilst going to infinite separation do not affect the pointed limit). One may view the local bounds as a translation of the usual criteria needed for Gromov’s precompactness theorem into more easily visualizable data on the size and spacing of subsets of metric spaces.
3.3 Tangent cones at infinity
@@
Finally, we show that for any finite dimensional normed vector space, there exists a smocked space whose unique tangent cone at infinity is that space. This answers in the affirmative an open question of Sormani et al. In general, there are examples of smocked spaces which do not have unique tangent cones at infinity, nor is the cone always a normed space [2].
Proposition 3.8: Let $(\mathbb{R}^{n},F)$ be a finite-dimensional normed vector space. Then there exists a smocked space $(X,d_{X})$ constructed in $\mathbb{E}^{n}$ and a basepoint $x_{0}\in X$ such that the rescaled pointed spaces $(X,\lambda^{-1}d_{X},x_{0})$ converge in the pointed Gromov-Hausdorff sense to $(\mathbb{R}^{n},F)$ as $\lambda\to\infty$. In particular, $(\mathbb{R}^{n},F)$ is the unique tangent cone at infinity of $(X,d_{X})$.
Proof: The tangent cone can be constructed explicitly using classical results on stable norms and homogenization [1, 7]. The first step is a standard one of approximating the ball by a rational polytope. Let $B_{F}=\{x:F(x)\leq 1\}$ be the unit ball of $F$. Since $B_{F}$ is convex and compact, there exists a symmetric finite subset $V=\{v_{1},\dots,v_{m}\}\subset\mathbb{Z}^{n}$ and weights $\ell_{i}:=F(v_{i})>0$ such that the polyhedral norm
$F_{V}(x):=\inf\Big\{\sum|a_{i}|\ell_{i}:\,x=\sum a_{i}v_{i}\Big\}$
satisfies $(1-\varepsilon)F(x)\leq F_{V}(x)\leq(1+\varepsilon)F(x)$ for all $x\in\mathbb{R}^{n}$ for arbitrarily small $\varepsilon>0$.
Let $L=\mathbb{Z}^{n}$ and define a graph $G=(V(G),E(G))$ with vertex set $V(G)=L$ and edges joining $p\in L$ to $p+v_{i}$ for every generator $v_{i}\in V$. Assign to the oriented edge $(p,p+v_{i})$ the length $\ell_{i}=F(v_{i})$. The induced shortest-path metric $d_{G}$ on $L$ is the weighted word metric associated to $V$. It is well known that the rescaled lattices $(L,\lambda^{-1}d_{G})$ converge in the pointed GH sense as $\lambda\to\infty$ to $(\mathbb{R}^{n},F_{V})$, and hence to $(\mathbb{R}^{n},F)$ because $F_{V}$ approximates $F$ uniformly.
Next, we must periodically embed this graph as a smocked subset of $\mathbb{E}^{n}$. This can be done as follows. For each vertex $p\in L$, insert a small closed ball $N_{p}:=\overline{B_{r}(p)}$ with radius $r>0$ chosen so that the balls (corresponding to nodes) are pairwise disjoint. For each edge $(p,p+v_{i})$, insert a narrow rectangular tube of Euclidean centerline length $\ell_{i}$ joining the boundaries $\partial N_{p}$ and $\partial N_{p+v_{i}}$. The tubes can be chosen to be sufficiently thin that distinct ones meet only at their endpoints. Let $S=\bigcup_{p\in L}N_{p}$ and define the smocked space $X:=\mathbb{E}^{n}/\!\!\sim$ by collapsing each node $N_{p}$ to a point via the smocking map $\pi:\mathbb{E}^{n}\to X$. This is a well-defined smocking of $\mathbb{E}^{n}$.
In the smocked space $X$, let $x_{p}=\pi(N_{p})$ for each vertex $p\in L$. For any two vertices $p,q\in L$, the distance between their images $x_{p}$ and $x_{q}$ equals the infimum of lengths of paths along connected tubes linking $N_{p}$ to $N_{q}$, since tubes are disjoint apart from their endpoints. This metric coincides with the weighted graph distance $d_{G}(p,q)$ up to an arbitrarily small error which can be eliminated by choosing the tube widths to be sufficiently small. It follows that the discrete subset $\{x_{p}\}_{p\in L}\subset X$ with induced metric $d_{X}$ is isometric to $(L,d_{G})$.
Since the construction is LL–periodic, the rescaled quotient metrics $(X,\lambda^{-1}d_{X})$ are periodic rescalings of the embedded weighted network. By the stable norm convergence and the fact that tube width can be neglected, it follows that
$(X,\lambda^{-1}d_{X},x_{0})\xrightarrow[\lambda\to\infty]{\mathrm{pGH}}(\mathbb{R}^{n},F),$
@@ 4.1 Measured convergence of smocked measure spaces
We now extend the preceding convergence results to smocked spaces endowed with measures. A natural choice of measure on a smocked space $X=\mathbb{E}^{N}/\!\!\sim$ is the pushforward of Lebesgue measure under the smocking map $\pi:\mathbb{E}^{N}\to X$,
$\mu:=\pi_{\#}(\mathcal{L}^{N}),$
which gives the total Euclidean volume of preimages in $\mathbb{E}^{N}$. This defines a canonical metric measure structure $(X,d_{X},\mu)$ associated to every smocked space. We show below that the pointed Gromov–Hausdorff convergence results from Section 3 extend naturally to this measured setting.
Proposition 4.1: Let $(X_{k},x_{k},d_{k})$ be smocked spaces obtained from $\mathbb{E}^{N}$ by collapsing collections of compact stitches $\{S_{\alpha}^{(k)}\}$, and let $\pi_{k}:\mathbb{E}^{N}\to X_{k}$ be the associated smocking maps. Assume that $(X_{k},x_{k})\xrightarrow{\mathrm{pGH}}(X_{\infty},x_{\infty})$, where $X_{\infty}$ is obtained from the limiting stitches $\{S_{\beta}^{(\infty)}\}$. Define measures
$\mu_{k}:=(\pi_{k})_{\#}(\mathcal{L}^{N}),\qquad\mu_{\infty}:=(\pi_{\infty})_{\#}(\mathcal{L}^{N}).$
Then
$(X_{k},d_{k},\mu_{k},x_{k})\xrightarrow{\mathrm{pmGH}}(X_{\infty},d_{\infty},\mu_{\infty},x_{\infty})\quad\text{as }k\to\infty.$
Proof: By Proposition 3.5, for every $R>0$ there exist $\varepsilon_{k}\to 0$ and Borel $\varepsilon_{k}$–isometries $f_{k}:B_{R}^{X_{k}}(x_{k})\to B_{R+\varepsilon_{k}}^{X_{\infty}}(x_{\infty})$ with $f_{k}(x_{k})=x_{\infty}$ and $B_{R-\varepsilon_{k}}^{X_{\infty}}(x_{\infty})\subset(f_{k}(B_{R}^{X_{k}}(x_{k})))_{\varepsilon_{k}}$. Since the maps $\pi_{k}$ and $\pi_{\infty}$ differ from the identity on $\mathbb{E}^{N}$ only within unions of small stitches whose total Euclidean measure tends to zero, the pushforward measures $\mu_{k}$ converge weakly to $\mu_{\infty}$ on bounded sets. Equivalently, for every bounded continuous $\varphi$ on $B_{R}^{X_{\infty}}(x_{\infty})$,
$\int_{X_{k}}\!\varphi\!\circ f_{k}\,d\mu_{k}=\int_{\mathbb{E}^{N}}\!\varphi(f_{k}\!\circ\!\pi_{k})\,d\mathcal{L}^{N}\longrightarrow\int_{\mathbb{E}^{N}}\!\varphi(\pi_{\infty})\,d\mathcal{L}^{N}=\int_{X_{\infty}}\!\varphi\,d\mu_{\infty}.$
@@
Corollary: If a sequence of smocking sets $S_{k}\subset\mathbb{E}^{N}$ converges in Hausdorff distance to $S_{\infty}$, and $(X_{k},d_{k})$, $(X_{\infty},d_{\infty})$ are the corresponding smocked spaces equipped with measures $\mu_{k}=(\pi_{k})_{\#}\mathcal{L}^{N}$ and $\mu_{\infty}=(\pi_{\infty})_{\#}\mathcal{L}^{N}$, then
$(X_{k},d_{k},\mu_{k},x_{k})\xrightarrow{\mathrm{pmGH}}(X_{\infty},d_{\infty},\mu_{\infty},x_{\infty}).$
This result shows that the smocking construction is stable under measured Gromov–Hausdorff convergence when equipped with the natural pushforward of Euclidean volume. In particular, the metric and measure structures both homogenize under the same geometric limits, confirming that smocked spaces extend consistently to the metric measure setting.
5 Conclusions
Smocked spaces, introduced by Sormani et al. as a geometric model for “stitched” or “collapsed” Euclidean structures, form a class of quotient spaces whose analytic and geometric properties have mostly been unexplored. In this work, we have provided a systematic treatment of their convergence theory and answered several open questions regarding their convergence properties. In particular, we have:
  • established the conditions under which Hausdorff convergence of the smocking sets implies pointed Gromov-Hausdorff convergence of the associated smocked spaces
  • identified local uniform bounds on the smocking constants as sufficient and necessary for Gromov–Hausdorff precompactness
  • constructed examples showing that global or uniform bounds are not required, thereby clarifying the geometric mechanisms underlying previous examples of convergence
  • proved that every finite-dimensional normed vector space arises as a tangent cone at infinity of a smocked space
  • extended the convergence theory to the setting of smocked metric measure spaces, showing stability under pointed measured Gromov–Hausdorff convergence.
These results shows that smocked spaces have a rigorous convergence theory similar to other established quotient space constructions in metric geometry, such as Alexandrov spaces and length-space gluings. The ’smocking’ concept provides a flexible mechanism for creating non-trivial limits of Euclidean spaces via controlled local collapse and hence these spaces may be a useful model for testing questions on metric convergence and tangent cones in a very geometrically explicit setting. They may also be useful as simplified models for metric collapse, where analytic and geometric quantities interact but can still be computed using elementary means.
The results developed here suggest several directions for future research. These include studying curvature or rectifiability properties of smocked spaces, formulating analogues of RCD-type or synthetic curvature conditions, and developing a corresponding theory of Gromov–Wasserstein convergence. In summary, this work demonstrates that smocked spaces are a non-trivial construction. Under appropriate geometric control, they exhibit rich convergence behavior and capture a wide variety of limit geometries. It would be interesting to investigate whether the smocking construction extends naturally to Wasserstein or Gromov–Wasserstein spaces, and whether analogues of contractivity or curvature dimension conditions hold for such smocked Wasserstein metrics.
Acknowledgments
This research was partly conducted whilst the author was visiting the Okinawa Institute of Science and Technology (OIST) through the Theoretical Sciences Visiting Program (TSVP).
Funding
This study was not funded.
Conflict of Interest
The author declares that he has no conflict of interest.
Ethical approval
This article does not contain any studies with human participants or animals performed by any of the authors.